OpenStudy (anonymous):

Differential Equations: Use method of reduction to find the second solution. t^2y''-2ty'-2y=0 y1(t)=t t>0

OpenStudy (anonymous):

\[t^2y''-2ty'-2y=0\\ y''-\frac{2}{t}y'-\frac{2}{t^2}y=0\] Let \(y_2(t)=v(t)y_1(t)=v(t)\cdot t,\) giving you \[y_2'(t)=t~v'(t)+v(t) \\ y_2''(t)=t~v''(t)+2v'(t)\] Substituting into the original equation, \[t~v''(t)+2v'(t)-\frac{2}{t}(t~v'(t)+v(t))-\frac{2}{t^2}t~v(t)=0\\ t~v''(t)+2v'(t)-2v'(t)-\frac{2}{t}v(t)-\frac{2}{t}~v(t)=0\\ t~v''(t)-\frac{4}{t}\color{red}{v(t)}=0\] The \(v(t)\) term should have disappeared (as is the case when doing reduction of order). I suspect you copied your diff.eq incorrectly.